OpenStudy (shubhamsrg):

limit n->infinity (1^k +2^k +3^k ....... n^k)/k(n^(k+1))

OpenStudy (experimentx):

could you clarify denominator??

OpenStudy (shubhamsrg):

\[\lim_{n \rightarrow infinity} (1^k +2^k ...... n^k)/(kn^{k+1})\]

OpenStudy (experimentx):

k times n^(k+1)??

OpenStudy (shubhamsrg):

yes..

OpenStudy (experimentx):

\[ \lim_{n \rightarrow \infty } {1 \over kn^{k + 1}}{\sum_{n=0}^\infty n^k}\]

OpenStudy (experimentx):

i guess we have to get the general formula for n^k

OpenStudy (shubhamsrg):

is there any?? i also thought of fancy things ,wolframed it,,but it said something about reiman zeta function.. but this is a simple 12th grade problem..

OpenStudy (shubhamsrg):

i somehow arrived at the ans by hit and trial by putting k=1,then 2 and 3 ..but thats cheating :|

OpenStudy (experimentx):

\[ \sum n^k = \sum (n+1)^{k}\] i'm not sure if we could use this technique .. looks lie long way to generalize it. can you post the wolf link?

OpenStudy (shubhamsrg):

maybe we can make that substitution, but any progress?

OpenStudy (shubhamsrg):

with that i mean **

OpenStudy (experimentx):

i guess that incorrect expression ... should be something like this http://www.wolframalpha.com/input/?i=+lim+n-%3Einf+sum%5Bi%5Ek%2C0%2Cn+%5D%2F%28kn%5E%28k%2B1%29%29 though wolf doesn't take it ... let me check it in mathematica first.

OpenStudy (experimentx):

Limit[Sum[i^k, {i, 0, n}]/(k n^(k + 1)), n -> Infinity] \[ = \text{Limit}\left[\frac{n^{-1-k} \left(0^k+\text{HarmonicNumber}[n,-k]\right)}{k},n\to \infty \right] \]

OpenStudy (anonymous):

\( \huge \lim_{n \rightarrow \infty} \frac{1^k+2^k+...+n^k}{k n^{k+1}}\\ \huge=\lim_{n \rightarrow \infty}\frac{1}{k} [(\frac{1}{n})^k+(\frac{2}{n})^k+...+(\frac{n}{n})^k] \frac{1}{n}\\ \huge \frac{1}{k} \int_{0}^{1} x^k dx=\frac{1}{k(k+1)} \)

OpenStudy (experimentx):

seems like i've been misinterpreting it whole time.

OpenStudy (anonymous):

i was wondering what u did that ;)

OpenStudy (experimentx):

reading 1^k + 2^k + 3^k ... as (1 + 2 + 3 + .. )^k lol

OpenStudy (anonymous):

lol :)

OpenStudy (experimentx):

anyway very nice and beautiful technique ... i haven't seen that before.

OpenStudy (experimentx):

i would never have never seen Riemann sums in that series.

OpenStudy (shubhamsrg):

excellent @mukushla ..thank you very much :)